Please show every step clearly! Also, The blue is the answers, I simply want to check my process.3. A steel tube (k = 15 W/mK) has an inner diameter of 40 mm and a wall thickness of 5 mm. The temperatures of the fluid insi
Answer #1

Soch gruen Ksteel = 15 w/mk ria 20mm Wall thickness - 5mm Ti=50 Ton 20°c hiz 300w/m2K ho=5w/milk. The heat gain per vuit lengdk 15 (0.02652)+(213-6762x103) +( 1*27323) CALC - = 11:52. Wing 11.52 Wlm 1.3024 hial un (8) 27 kstell ho to in (2) 27 KthrekD ove s9 +(23476-18) + 2. (0.02652)+( 2.3676x103) + 2.6975 + 0.9094 S - 4.15 w/m swim = 2.6118 Ay